A large retailer obtains merchandise under the credit terms of 1/10, net 45, but routinely takes 60 days to pay its bills. (Because the retailer is an important customer, suppliers allow the firm to stretch its credit terms.) What is the retailer's effective cost of trade credit? Assume 365 days in year for your calculations. Do not round intermediate calculations.
Answer #1

Step 1: Calculation of Periodic Rate

Here the credit terms of 1/10, net 45. That is the discount percentage is 1% if the payments are made within 10 days instead of 45 days.

Periodic Rate = Discount percentage/(1-Discount percentage)

= .01/(1-.01)

= .01/.99

= 0.0101

Step 2: Calculation of Periods per year

Periods per year = 365/(Days allowed-Discount period)

= 365/(60-10)

= 365/50

= 7.3

Step 3: Calculation of Effective Cost of Trade Credit

Effective Cost of Trade Credit = ((1+Periodic Rate)^Periods per year) - 1

= ((1+0.0101)^7.3) - 1

= 1.0761180329-1

= 0.0761180329

= 7.61%